how to factor
4y^2-14y+6

Answer by stanbon(75887)

how to factor
4y^2-14y+6
-------
= 2(2y^2-7y+3)
-------
= 2(2y^2-6y-y+3)
-------
= 2(2y(y-3)-(y-3))
-------
= 2(y-3)(2y-1)
-----------------
Cheers,
Stan H.
------